APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday December 29, 2009 (MJD 55195) ---=== OBSERVING TEAM ===--- Viktor Malanushenko Audrey Simmons Kaike Pan (afternoon) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- We did not open tonight because of falling snow (additional 1" or so) and high humidity. We completed dithered flats manually. There were a few times when boss moveColl was issued and the motor positions moved the correct amount, but error messages appeared in the log saying that the motor positions failed. These messages appeared immediately after the command was issued (we were moving the collimators by 62 ticks at a time). A flat field bulb burned out during the test. We replaced it. ---=== OBSERVING LOG ===--- Afternoon: --------- All actors and systems were checked and conformed working properly. High humidity and snowing prevents from ventilation. Shelved snow from walks. Night: ----- We could not open due to snow, clouds, and high humidity. *** We began gathering dithered flats. 00:57 sop hartmann 105661, 662 00:57 sp1 105661 16 3693 arc Lef 4.0 6913 7559 7388 HART 00:57 sp1 105661 16 3693 arc Lef 5.0 14854 14599 14680 HART 00:59 sp2 105662 16 3693 arc Rig 5.0 6913 7559 7388 HART 00:59 sp2 105662 16 3693 arc Rig 5.0 14854 14599 14680 HART 01:06 boss moveColl spec=sp1 a=2406 b=2406 c=2406 01:07 boss moveColl spec=sp2 a=2075 b=2075 c=2075 01:10 sop hartmann 105663, 664 - it to focus position 01:10 sp1 105663 16 3693 arc Lef 5.0 9304 9966 9792 HART 01:10 sp1 105663 16 3693 arc Lef 5.0 16926 16673 16755 HART 01:12 sp2 105664 16 3693 arc Rig 4.0 9304 9966 9792 HART 01:12 sp2 105664 16 3693 arc Rig 5.0 16926 16673 16755 HART COLLIMATE: Camera b1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of 3.1 degrees for spectro-1 COLLIMATE: Camera r1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of 244 steps for spectro-1 COLLIMATE: Camera b2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of -1.7 degrees for spectro-2 COLLIMATE: Camera r2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of 691 steps for spectro-2 01:29 sop doCalibs narc=1 -- arc 'in foc' position, 105665 01:29 105665 16 3693 arc Out 4.0 9304 9966 9792 16926 16673 16755 01:38 boss moveColl spec=sp1 a=682 b=682 c=-682 boss moveColl spec=sp2 a=682 b=682 c=-682 01:41 sop doCalibs narc=1 -- in to starting==1st position, 105666 01:41 105666 16 3693 arc Out 4.0 9986 10648 9112 17609 17354 16073 1st test of dithered flats: --------------------------- 1 11 - 15 01:48 105667 16 3693 fla Out 31.0 9986 10648 9112 17609 17354 16073 2 16 - 20 02:04 105672 16 3693 fla Out 31.0 9933 10586 9167 17547 17293 16145 3 21 - 25 02:19 105677 16 3693 fla Out 31.0 9864 10524 9233 17485 17231 16205 4 26 - 30 02:34 105682 16 3693 fla Out 31.0 9805 10462 9293 17423 17169 16270 5 31 - 35 02:49 105687 16 3693 fla Out 31.0 9743 10399 9356 17361 17107 16321 6 36 - 40 03:06 105692 16 3693 fla Out 31.0 9681 10338 9418 17299 17045 16383 7 41 - 45 03:22 105697 16 3693 fla Out 30.0 9621 10275 9480 17237 16983 16454 8 46 - 50 03:40 105702 16 3693 fla Out 30.0 9560 10213 9540 17175 16921 16514 9 51 - 55 03:54 105707 16 3693 fla Out 31.0 9498 10151 9603 17113 16860 16569 10 56 - 60 04:09 105712 16 3693 fla Out 31.0 9434 10090 9664 16989 16736 16693 11 61 - 65 04:25 105717 16 3693 fla Out 31.0 9373 10027 9724 16989 16736 16693 12 66 - 70 04:40 105722 16 3693 fla Out 31.0 9311 9965 9786 16927 16674 16763 13 71 - 75 04:57 105727 16 3693 fla Out 31.0 9247 9903 9850 16865 16612 16817 14 76 - 80 05:11 105732 16 3693 fla Out 31.0 9187 9841 9913 16803 16550 16879 15 81 - 85 05:25 105737 16 3693 fla Out 31.0 9127 9779 9968 16741 16488 16941 16 86 - 90 05:39 105742 16 3693 fla Out 31.0 9066 9717 10033 16679 16426 17011 17 91 - 95 05:54 105747 16 3693 fla Out 31.0 9000 9655 10095 16617 16364 17065 18 96 -100 06:09 105752 16 3693 fla Out 31.0 8940 9593 10157 16555 16302 17127 19 101 -105 06:23 105757 16 3693 fla Out 30.0 8878 9531 10219 16493 16240 17198 21 106 -110 06:39 105762 16 3693 fla Out 31.0 8754 9407 10346 16369 16116 17321 22 111 -115 06:53 105767 16 3693 fla Out 31.0 8694 9345 10406 16307 16054 17375 23 116 -120 07:08 105772 16 3693 fla Out 30.0 8629 9283 10466 16246 15992 17437 20 121 -124 07:26 105777 16 3693 fla Out 31.0 8805 9469 10290 16431 16178 17251 here is only 4 exposures (i.e. at last exposure one of the FF lamp did not "ON") 2nd test: --------- -- in collimate 07:52 105781 16 3693 fla Out 30.0 9307 9966 9792 16927 16673 16754 07:59 105782 16 3693 arc Out 5.0 9307 9966 9792 16927 16673 16754 08:01 105783 16 3693 fla Out 31.0 9307 9966 9792 16927 16673 16754 -- -651, -651, +651 : 08:14 105784 16 3693 arc Out 5.0 8656 9313 10438 16275 16021 17406 08:17 105785 16 3693 fla Out 31.0 8656 9313 10438 16275 16021 17406 -- +1302, +1302, -1302 : 08:58 105787 16 3693 fla Out 31.0 9958 10618 9140 17577 17323 16104 ----------------------- ======================= the motors come back in to 'in foc'= collimate position check collimations: sop hartmann: expositions 105788, 789: 09:23 105788 16 3693 arc Lef 5.0 9304 9973 9796 HART 09:23 105788 16 3693 arc Lef 5.0 16926 16673 16755 HART 09:26 105789 16 3693 arc Rig 5.0 9304 9973 9796 HART 09:26 105789 16 3693 arc Rig 5.0 16926 16673 16755 HART results: COLLIMATE: Camera b1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of 5.0 degrees for spectro-1 COLLIMATE: Camera r1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of -489 steps for spectro-1 COLLIMATE: Camera b2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of -0.6 degrees for spectro-2 COLLIMATE: Camera r2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of -184 steps for spectro-2 Troubleshooting: ================ -->> SYMPTOM: scientific exposure can not reading, and one time per min 'boss' try 'prereading': *** .boss 7371 boss w text="exposureState timer does not add up (remain=-317.60 total=3.00)" *** .boss 7371 boss i exposureState=PREREADING,3.0,0.0 exposure status: log> boss exposure status 10:32:12 130 boss exposure status 10:32:12 .hub 0 cmds d CmdIn="APO.elenam","boss","exposure status" 10:32:12 .hub 0 cmds d CmdQueued=104695,1261305133.32,"APO.elenam",130,"boss",216,"exposure status" 10:32:12 APO.viktorm.sop 279 boss w text="exposureState timer does not add up (remain=-667.27 total=3.00)" 10:32:12 APO.viktorm.sop 279 boss i exposureState=PREREADING,3.0,0.0 SOLUTION: log> boss exposure abort log> boss exposure readout 10:41Z boss exposure abort 10:42Z boss exposure readout <<-- -------------------- -->> SYMPTOM: you run command 'sop doScience...' and no answer in to bossGUI window SOLUTION: log> sop restart and against log> sop doScience ... -->> SYMPTOM: the 'mcp' does not worked or worked strangely=non-standart SOLUTION: log> mcp info 08:35Z - FF bulb burned out (at 3-6:00 position). Had to replace it. 09:20Z - Finished dithered flats. *** 11:07Z - endNight finished. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:24Z 23 20 0 210 (SSW) 0 - 22:56Z 27 21 9 326 (NW) 0 - 23:28Z 24 21 12 304 (NW) 0 - 00:00Z 25 21 16 293 (WNW) 0 - 00:32Z 25 21 20 298 (WNW) 0 - 00:57Z 25 21 16 294 (WNW) 0 - 01:29Z 24 21 16 289 (WNW) 0 - 02:01Z 24 21 21 294 (WNW) 0 - 02:33Z 23 20 11 308 (NW) 0 - 03:05Z 24 21 18 288 (WNW) 0 - 03:37Z 23 20 13 296 (WNW) 0 - 04:09Z 24 20 14 307 (NW) 0 - 04:42Z 24 21 16 274 (W) 0 - 05:14Z 24 21 18 308 (NW) 0 - 05:46Z 24 21 23 269 (W) 0 - 06:18Z 24 21 25 265 (W) 0 - 06:50Z 24 21 22 260 (W) 0 - 07:22Z 24 21 22 271 (W) 0 - 07:54Z 24 21 29 287 (WNW) 0 - 08:27Z 24 21 23 263 (W) 0 - 08:59Z 24 21 20 281 (W) 0 - 09:31Z 25 21 30 267 (W) 0 - 10:03Z 25 21 30 266 (W) 0 - 10:35Z 25 21 23 269 (W) 0 - 11:08Z 25 21 20 278 (W) 0 - ---=== TELESCOPE STATUS ===--- 10:53Z - Enclosure is on, doors and louvers are closed, and fans are off. No interlocks bypassed. Status at: 10:53Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 6 Counterweights at: 250 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 175.30 lbs, 20.80 psi Imager lbs, psi (IMAGER IS WARM) ---=== SOFTWARE USED ===--- IOP/SOP: v4_40_0 Watcher: v2_36_0 MCP: v6_6_1 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.11.0 2009-06-24 sdssProcedures: SoS: plate-mapper: ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.7040 -0.6940 Axial B -0.0250 -0.0410 Axial C -0.3790 -0.3710 Trans D 11.6120 11.6170 Lateral E 10.3759 10.3505 Lateral F 11.9507 11.9253 GALILS Commanded: 9900. 400. 5850. 1200. 2950. 2750. Actual: 9900. 400. 5850. 1200. 2950. 2750. SETMIR VALUES PriDesOrient: 0.00 -3.30 5.10 278.08 178.06 PriOrient: 0.00 -3.04 5.26 277.70 178.07 SECONDARY: ---------- Focus: 0.00 Air Temp.: -3.8 Alt.: 29.999997 MIGS TONIGHT NOMINAL Axial A 18.6150 18.6150 Axial B 15.6280 15.6120 Axial C 13.8270 13.8170 Trans D -11.7750 -11.7780 GALILS Commanded: 1600342. 1564537. 1585854. -7600. -12400. Actual: 1597700. 1566200. 1588100. -7600. -12400. SETMIR VALUES SecDesOrient: 1257.00 -4.00 -9.00 10.00 234.42 SecOrient: 1256.66 -3.31 -10.46 9.23 234.24 ---=== PROBLEMS IN DETAIL ===--- .... 02:21:47 APO.viktorm.sop 960 boss w text="failed to make SPA card from None" ??? ... 03:04:35 APO.viktorm 33 boss f text="Motor failed to move motor c on sp1mech" ... 03:38:12 .alerts 0 alerts i alert=imager.heartbeat,critical,"The imager's been dead for at least 183314s",disabled,noack,"" 03:38:13 APO.viktorm 39 boss w text="Failed message from sp1mech, failed: { motor timeout }" 03:38:13 APO.viktorm 39 boss f text="Motor failed to move motor c on sp1mech" 03:38:13 .hub 0 cmds d CmdDone=184344 ... 03:38:14 .boss 945 boss w text="this command has already been finished!!!! (APO.viktorm 945): moveColl spec=sp1 a=-62 b=-62 c=62" 03:38:14 .boss 945 boss d text="Completed message from sp1mech" 03:38:14 .boss 945 boss w text="this command has already been finished!!!! (APO.viktorm 945): moveColl spec=sp1 a=-62 b=-62 c=62"